I have to solve each system of equations by the elimination method.
3/4x + 1/3y = 1
1/2x + 2/3y = 0

Answer by checkley77(12844)

3x/4+y/3=1 multiply this equation by -2 & add them.
x/2+2y/3=0
-6x/4-2y/3=-2
--------------------
x/2-6x/4=-2
(2x-6x)/4=-2
-4x/4=-2
-4x=-8
x=-8/-4
x=2 ans.
2/2+2y/3=0
1+2y/3=0
2y/3=-1
2y=-3
y=-3/2 ans.
Proof:
3*2/4+(-3/2)/3=1
6/4-3/2*1/3=1
6/4-3/6=1
1.5-.5=1
1=1